# Sequence Generation Modeling for Continuous Value Prediction

### v2

@@ 3.4. Label Encoding
Given the vocabulary $\mathcal{V}$, we need to do label encoding, i.e., transforming the watch time values $\{y_{i}\}_{i=1}^{N}$ to corresponding target sequences $\{\bm{s_{i}}=\{s_{i}^{1},...,s_{i}^{T_{i}}\}\}_{i=1}^{N}$. To facilitate the learning process with the encoded labels (i.e., transformed token sequences), we also offer three principles for label encoding as follows:
  • Correctness: $y_{i}=\sum_{t=0}^{T_{i}}w_{i}^{t}+\epsilon$, where $\epsilon$ is a small value (e.g. 0.1% relative error) which can be omitted.
  • Minimal sequence length: $T_{i}$ is the possible minimal length of the token sequence selected to satisfy correctness for $y_{i}$.
  • Monotonicity: $s_{i}^{1}\geq s_{i}^{2}\geq...\geq s_{i}^{T_{i}}$ is required.
The minimum sequence length requirement simplifies the learning process. Additionally, the non-decreasing order captures the dynamic changes in user interest while watching videos, making the sequence generation process progressively (from easy to hard). Without loss of generality, the token indices in the vocabulary are also sorted in descending numerical order.
To this end, we employ a greedy strategy for the shortest sequence length, starting from the largest possible watch time slot and decreasing progressively, decomposing the total watch time $y_{i}$ into a sequence of watch time slots.
@@ 3.5.1. Vanilla Training Process
Similar to language modeling, the model predicts the next token $s^{t}$ using the ground truth tokens $s^{<t}$. The learning objective is to maximize the negative log-likelihood of the generated sequence over the training set and effectively minimize the difference between the predicted sequence and the ground truth sequence, which is computed by the cross-entropy loss as follows:
@@
where $\hat{y_{i}}=\sum_{t=1}^{T_{i}}g(\hat{s}_{i}^{t})$, $\delta$ acts as a threshold, toggling between quadratic and linear losses to balance sensitivity and robustness against outliers. Therefore, the loss function we need to optimize is
@@
$\lambda$ is a hyperparameter that balances the two losses. To improve model efficiency, we adopt a teacher forcing (TF) strategy, which directly feeds the ground truth output $s_{i}^{t}$ as input at step $t+1$ to guide model training. However, since the ground truth is unknown during inference, the discrepancy in input for the decoder leads to the well-known exposure bias problem, which can degrade model performance (Goodman et al., 2020).
@@ 3.5.2. Curriculum Learning with Embedding Mixup (CLEM)
We introduce a Curriculum Learning (CL) strategy that combines TF and vanilla training (a.k.a free running mode) by progressively incorporating stochasticity into the training process. Specifically, to predict $\hat{s}_{i}^{t}$, we randomly choose ground truth token $s_{i}^{t-1}$ with a dynamic probability $p$ as the sampling rate.
@@
We notice that GR can capture the inherent relationship between different tokens. As the cardinality of $\mathcal{V}$ is not as large as that in LMs, it is hard to visualize $\mathcal{V}$’s embedding matrix $\bm{E}$. As an alternative, we visualize the target numerical embeddings of $y_{i}$ by $\bm{e_{i}}=\sum_{t=1}^{T_{i}}r_{t}\bm{E}[\hat{s_{i}}^{t},:]$, $r_{t}=\frac{g(\hat{s}_{i}^{t})}{y_{i}}$ is the contribution of $\hat{s}_{i}^{t}$ to the target value $y_{i}$. As shown in Fig. 3(a), values beginning with the same token $w_{j}$ are clustered, and the initial tokens of similar numeric values $\text{g}(w_{j})$ are also relatively close to each other. This clustering feature indicates a coherent structure in the embedding space, suggesting that the model effectively captures similarities among the tokens.
Although this is what we expect the model to learn, it can also lead to the issue of confusion among similar tokens. In other words, the model may struggle to effectively distinguish between the target token and its neighboring tokens. As noted in Sec. 3.4, tokens are arranged in non-increasing order within the vocabulary $\text{g}(w_{1})>\text{g}(w_{2})>...>\text{g}(w_{V})$. We illustrate this phenomenon in Fig. 3(b) by computing the averaged probability difference of each token relative to its neighbors. The results indicate that the predicted probability of the target token is closely aligned with the surrounding tokens, which contributes to an increase in the final prediction error. Since our token sequences are encoded in non-increasing order, it becomes challenging for subsequent tokens to bridge the gap between $\hat{y}_{i}$ and ground truth $y_{i}$ once the first token, denoted as $\hat{s_{i}}^{1}$ is wrong. Additionally, the argmax operation in Eq. (2) will disrupt the gradient propagation path, which further complicates the learning process.
We observe that employing a curriculum training strategy with inference by using a beam search ensemble significantly boosts model performance. However, it incurs more computational cost. Given the previously discussed properties of tokens and sequences, the numerical differences between adjacent tokens are minimal. To address this, we propose to integrate the embedding sequences of the preceding tokens through an ensemble approach during the training process to improve the prediction of the next token.
@@
Embedding Mixup (EM). The cohort is centered on the current predicted token $\hat{s}^{t}_{i}$ with window size $n_{w}$, the mixup region is $[\hat{s}^{t}_{i}-b,\hat{s}^{t}_{i}+b]$ and $b=\lfloor\frac{n_{w}}{2}\rfloor$. We have
(8) $\bm{z_{i}}^{t}=\sum\limits_{j=0}^{n_{w}}\sigma_{j}\cdot\bm{E}[\hat{s}^{t}_{i}+j-b,:]$
(9) $\sigma_{j}=\frac{exp(-\rho_{j})}{\sum_{k=0}^{n_{w}}exp(\rho_{\hat{s}^{t}_{i}+k-b})}$
@@
where $\sigma_{j}$ recalculates the fusion weights of tokens in the fixed window size, $\rho_{j}$ is the logit predicted by the decoder at step $t$. And $\bm{z_{i}}^{t}$ will replace original $\bm{E}[\hat{s}^{t}_{i},:]$ as the input for the next step. This re-weighted mixup approach leverages the numerical semantics and additivity property inherent in our tokens. The re-weighting ensures that the semantic space is aligned, eliminating any discrepancies in scale. EM offers three benefits: 1) It reduces the learning complexity of the model by merging representations of tokens within a fixed window, thereby preventing significant errors; 2) The integration leverages the predicted scores from the previous steps, enhancing the information transfer from output to input in recurrence structure; and 3) it ensures consistency between training and inference while lowering inference cost.
@@ 3.5.3. Inference Process
During inference, the encoder extracts $\bm{h_{i}}$ from input features $[\bm{u_{i}},\bm{v_{i}}]$, the decoder begins with the ¡SOS¿ token, and sequentially generates the prediction sequence $\hat{s}_{i}=\{\hat{s}_{i}^{1},\hat{s}_{i}^{2},...,\hat{s}_{i}^{T_{i}}\}$. The process continues until the token ¡EOS¿ is generated, which signifies the completion of the sequence. Finally, the predicted watch time is computed as $\hat{y_{i}}=\sum_{t=1}^{T_{i}}\text{g}(\hat{s}_{i}^{t})$.
@@ 4. Experiments
  • RQ2: What is the effect of vocabulary design on the performance of GR and why?
  • RQ3: How do various neural network architectures perform within the GR framework?
@@
Due to the space limit, additional ablation study results are available at A.2.
@@ 4.1. Experiment Settings
4.1.1. Datasets
Two publicly available datasets and one industrial dataset are used to evaluate the proposed method. The large-scale industrial dataset (Indust for short) is sourced from a real-world streaming short-video app Kuaishou with about 400 million DAUs and multi-billion impressions each day. We collect interaction logs for 4 days and utilize the subsequent day’s data for evaluation. The CIKM16333https://competitions.codalab.org/competitions/11161, sourced from the CIKM16 Cup competition, is designed to predict user engagement duration in online search sessions. It contains 310,302 sessions, 122,991 items, and an average session length of 3.981. KuaiRec (Gao et al., 2022a) is a real-world dataset collected from the video view logs of the Kuaishou app, comprising data from 7,176 users, 10,728 items, and 12,530,806 impressions. In addition to directly predicting watch time, some researchers (Lin et al., 2023) focus on predicting the watch ratio, which can be used in conjunction with video duration to calculate watch time. Therefore, we report both watch time and watch ratio results on the KuaiRec dataset.
Table 2. Performance comparison among different approaches on KuaiRec, CIKM16 and Indust.
@@ 4.1.1. Datasets
VR 7.634 0.504 - 0.385 0.691 - 1.039 0.641 - 46.343 0.588
WLR (Covington et al., 2016) 6.047 0.525 4.167% 0.375 0.698 1.013% 0.998 0.672 4.836% - -
D2Q (Zhan et al., 2022) 5.426 0.565 12.103% 0.371 0.712 3.039% 0.899 0.661 3.120% - -
TPM (Lin et al., 2023) 3.456 0.571 13.294% 0.361 0.734 6.223% 0.850 0.676 5.460% 41.486 0.593
CREAD (Sun et al., 2024) 3.307 0.594 17.857% 0.369 0.738 6.802% 0.865 0.678 5.772% 39.979 0.597
GR (ours) 3.196 0.614 21.825% 0.333 0.753 8.972% 0.815 0.691 7.80% 38.528 0.604
@@
A/B test APP usage time Average app usage per user
+0.055% +0.072%
APP launch counts Video consumption time
+0.042% +0.129%
4.1.2. Compared Methods
Considering baseline methods compared in prior studies (Sun et al., 2024; Lin et al., 2023), we compare several state-of-the-art methods (Covington et al., 2016; Zhan et al., 2022; Sun et al., 2024; Lin et al., 2023) with our GR. For a fair comparison, similar model structures except for the output layers and corresponding loss functions are employed for these approaches. Details of the compared methods are provided in ”watch time prediction” of Sec. 2.
4.1.3. Metrics
  • Mean Average Error (MAE): It quantifies regression accuracy by averaging the absolute deviations between predicted values $\{\hat{y_{i}}\}_{i=1}^{N}$ and actual values $\{y_{i}\}_{i=1}^{N}$ as follows:
    (10) $MAE=\frac{1}{N}\sum_{i=1}^{N}\left|\hat{y}_{i}-y_{i}\right|$
4.2. Performance Comparison (RQ1)
4.2.1. Offline
We compare the proposed method GR with 5 existing models in terms of watch time and watch ratio, results are presented in Tab. 2. Specifically, for watch time prediction, GR exhibits a 4.117% relative decrease in MAE and a 1.917% improvement in XAUC on the CIKM16 dataset. On the KuaiRec benchmark, significant improvements are achieved, with a 3.356% reduction in MAE and a 3.367% increase in XAUC relative to the second-best method. As for Indust dataset, GR exhibits a 3.629% relative decrease in MAE and a 1.001% improvement in XAUC compared to the CREAD. Regarding watch ratio predictions, all models exhibit significant performance gain due to the elimination of video duration bias. Despite this, GR maintains the best performance, boasting a 7.756% reduction in MAE and a 2.033% improvement in XAUC. The comprehensive improvements in both MAE and XAUC substantiate the superiority of the GR method. To confirm that the observed enhancements in the GR model are not solely attributable to increased model parameters, further comparisons with parameter-equivalent models can be found in the appendix A.2.1.
4.2.2. Online A/B Testing
Besides offline experiments, we also conduct an online A/B test on Kuaishou App to demonstrate the real-world efficacy of our method. Considering that Kuaishou serves about 400 million users daily, doing experiments from 2.5% of traffic involves a huge population of more than 10 million users, which can yield highly reliable results. Most recommendation systems follow a two-stage framework: initially retrieving a set of candidate items and then selecting the top-ranked items in the ranking phase. Watch time prediction, a component of the ranking stage, plays a crucial role. Items with higher predicted watch times are more likely to be recommended.
The online experiment have been launched on the system for six days and the results are in Tab. 3. The metrics for online experiments include app usage time, average app usage per user, app launch counts and video consumption time (accumulated watch time). The results demonstrate that GR consistently boosts performance in watch time related metrics, with an improvement by 0.055% on app usage time, 0.072% on average app usage per user, and significant 0.129% on video consumption time, substantiating its potential to significantly enhance real-world user experiences.
4.3. Vocabulary Construction Analysis (RQ2)
@@
[…]
@@
4.4. Generative Architecture Analysis (RQ3)
To investigate the impact of different network architectures on prediction performance, we compare three popular sequence modeling structures: RNN, GRU, and LSTM, by substituting the decoder structure in the GR model. Results are presented in Fig. 5, from which we can see that GR is not dependent on a specific network architecture and achieves desirable outcomes across various architectures. Among them, LSTM exhibits the best performance, likely because GR ultimately sums all predicted tokens to estimate watch time, resulting in strong dependencies between tokens in the sequence. LSTM is particularly adept at capturing these dependencies across different token positions.

### v3

@@ 3.4. Label Encoding
Given the vocabulary $\mathcal{V}=\{w_{1},w_{2},...,w_{V}\}$, we perform label encoding to transform the watch time values $\{y_{i}\}_{i=1}^{N}$ into corresponding target sequences $\{\bm{s}_{i}=\{s_{i}^{1},\ldots,s_{i}^{T_{i}}\}\}_{i=1}^{N}$. To guide the label encoding process, we propose three foundational principles:
  • Correctness: The original value must be reconstructible from the token sequence with bounded error:
    (4) $y_{i}=\sum_{t=1}^{T_{i}}g(s_{i}^{t})+\epsilon,\quad\text{where }|\epsilon|\leq 0.001\cdot y_{i}$
  • Minimal Sequence Length: The sequence length $T_{i}$ should achieve the minimal possible cardinality while satisfying the correctness constraint.
  • Monotonicity: Token values must satisfy a non-increasing order:
    (5) $g(s_{i}^{1})\geq g(s_{i}^{2})\geq\cdots\geq g(s_{i}^{T_{i}})$
The minimum sequence length principle reduces learning complexity, while the monotonic constraint captures decaying user attention patterns during video watching.
To follow these principles, we implement a greedy decomposition algorithm. Starting from the largest possible watch time slot and decreasing progressively, decomposing the total watch time $y_{i}$ into a sequence of watch time slots.
@@ 3.5.1. Vanilla Training Process
Following language modeling paradigms, the model predicts the next token $s^{t}$ conditioned on preceding ground truth tokens $s^{<t}$. The learning objective minimizes the cross-entropy loss between predicted and ground truth sequences:
@@
where $\hat{y_{i}}=\sum_{t=1}^{T_{i}}g(\hat{s}_{i}^{t})$, $\delta$ acts as a threshold, toggling between quadratic and linear losses to balance sensitivity and robustness against outliers. Therefore, the composite loss becomes:
@@
where $\lambda$ is a hyperparameter that balances the two losses. To improve model efficiency, we adopt a teacher forcing (TF) strategy (Venkatraman et al., 2015), which directly feeds the ground truth output $s_{i}^{t}$ as input at step $t+1$ to guide model training. However, since the ground truth is unknown during inference, the discrepancy of input for the decoder leads to the well-known exposure bias problem (Goodman et al., 2020), which can degrade model performance.
@@ 3.5.2. Curriculum Learning with Embedding Mixup (CLEM)
To mitigate exposure bias inherent in teacher forcing, we propose a phased Curriculum Learning (CL) strategy. Specifically, to predict $\hat{s}_{i}^{t}$, we randomly choose ground truth tokens $s_{i}^{t-1}$ or predicted tokens $\hat{s}_{i}^{t-1}$ with a dynamic probability $p$ as the sampling rate. However, Transformer processes the entire sequence in parallel during a single forward pass, preventing access to the predicted tokens of previous time steps. Thus, as shown in Fig. 2, we implement CL with two forward passes through the decoder during training. The first pass performs vanilla training to obtain initial model predictions. In the second pass, inputs are sampled between ground truth tokens and predicted tokens with probability $p$, yielding the final predictions. Both passes share the same model parameters.
@@
Our analysis reveals that GR effectively captures inter-token relationships through its embedding structure. Given the vocabulary size $V$ being orders of magnitude smaller than typical language models, we analyze token semantics via aggregated value embeddings:
(10) $\bm{e}_{i}=\sum_{t=1}^{T_{i}}r_{t}\bm{E}[s_{i}^{t},:],\quad r_{t}=\frac{g(s_{i}^{t})}{y_{i}}$
where $r_{t}$ represents the contribution of token $s_{i}^{t}$ to target value $y_{i}$. Fig. 3(a) demonstrates two key properties:
  • Token Clustering: Values sharing initial tokens form distinct clusters.
  • Semantic Continuity: Embeddings of tokens with similar $g(w_{j})$ values reside in proximate regions.
This structural coherence facilitates numerical reasoning through geometrically meaningful representations. As noted in Sec. 3.4, tokens are arranged in non-increasing order within the vocabulary $\text{g}(w_{1})>\text{g}(w_{2})>...>\text{g}(w_{V})$. We also compute the averaged probability difference of each token relative to its neighbors and observe that tokens with neighboring indices in the vocabulary demonstrate the highest probability similarity in the model’s predictions, as shown in Fig. 3(b).
@@
To improve the prediction precision of next token, we propose to integrate the embedding sequences of the preceding tokens through a local ensemble approach called Embedding Mixup (EM) during the training process. The cohort is centered on the current predicted token $\hat{s}^{t}_{i}$ with window size $n_{w}$, the mixup region is $[\delta_{\hat{s}^{t}_{i}}-b,\delta_{\hat{s}^{t}_{i}}+b]$ and $b=\lfloor\frac{n_{w}}{2}\rfloor$, $\delta_{\hat{s}^{t}_{i}}$ represents the token index of $\hat{s}^{t}_{i}$ in $\mathcal{V}$. We have
(11) $\bm{z_{i}}^{t}=\sum\limits_{j=0}^{n_{w}}\sigma_{j}\cdot\bm{E}[\delta_{\hat{s}^{t}_{i}}+j-b,:]$
(12) $\sigma_{j}=\frac{exp(-\rho_{j})}{\sum_{k=0}^{n_{w}}exp(\rho_{\delta_{\hat{s}^{t}_{i}}+k-b})}$
@@
where $\bm{E}\in\mathbb{R}^{V\times D}$ is the vocabulary embedding matrix, $\sigma_{j}$ recalculates the fusion weights of tokens in the fixed window size, $\rho_{j}$ is the logit predicted by the decoder at step $t$. Specifically, $\bm{z_{i}}^{t}$ will replace original $\bm{E}[\hat{s}^{t}_{i},:]$ as the input at $t+1$ step. This re-weighted EM approach leverages the numerical semantics and additivity property inherent in our tokens. The re-weighting ensures that the semantic space is aligned, eliminating any discrepancies in scale. EM offers three benefits: 1) It reduces the learning complexity of the model by merging representations of tokens within a fixed window, thereby preventing significant errors; 2) The integration leverages the predicted scores from the previous steps, enhancing the information transfer from output to input in recurrence structure and restructuring the gradient propagation path; and 3) it ensures consistency between training and inference while lowering inference cost.
@@ 3.5.3. Inference Process
During inference, the encoder extracts $\bm{h_{i}}$ from input features $[\bm{u_{i}},\bm{v_{i}}]$, the decoder begins with the ¡SOS¿ token and sequentially generates the prediction sequence $\hat{s}_{i}=\{\hat{s}_{i}^{1},\hat{s}_{i}^{2},...,\hat{s}_{i}^{T_{i}}\}$, with each token generated using only the first forward pass. The process continues until the token ¡EOS¿ is generated, which signifies the completion of the sequence. Finally, the predicted watch time is computed as $\hat{y_{i}}=\sum_{t=1}^{T_{i}}\text{g}(\hat{s}_{i}^{t})$.
@@ 4. Experiments
  • RQ2: What are the underlying reasons behind the model’s performance exceeding the baseline?
  • RQ3: What is the effect of vocabulary design on the performance of GR and why?
@@ 4.1. Experiment Settings
4.1.1. Datasets.
We evaluate our method on one industrial dataset and two public benchmarks. The large-scale industrial dataset (Indust for short) is sourced from a real-world short-video app Kuaishou with over 400 million DAUs and multi-billion impressions each day. We use interaction logs spanning 4 days for training and those from the subsequent day for testing. We also use the public CIKM16333https://competitions.codalab.org/competitions/11161 and KuaiRec (Gao et al., 2022a) datasets, adopting the experimental settings from previous works (Lin et al., 2023; Sun et al., 2024) (Details are provided in the supplementary material). Consistent with prior work (Lin et al., 2023), we also report watch ratio results on KuaiRec, which can be used in conjunction with video duration to calculate watch time.
4.1.2. Metrics
  • Mean Average Error (MAE): It quantifies regression accuracy by averaging the absolute deviations between predicted values $\{\hat{y_{i}}\}_{i=1}^{N}$ and actual values $\{y_{i}\}_{i=1}^{N}$ by $\frac{1}{N}\sum_{i=1}^{N}\left|\hat{y}_{i}-y_{i}\right|$.
Table 1. Performance comparison among different approaches on KuaiRec, CIKM16 and Indust dataset.
@@ 4.1.2. Metrics
VR 7.634 0.534 - 0.385 0.691 - 1.039 0.641 - 46.343 0.588
WLR (Covington et al., 2016) 6.047 0.545 2.059% 0.375 0.698 1.013% 0.998 0.672 4.836% - -
D2Q (Zhan et al., 2022) 5.426 0.565 8.757% 0.371 0.712 3.039% 0.899 0.661 3.120% - -
CWM (Zhao et al., 2024) 3.452 0.580 8.614% 0.368 0.725 4.920% 0.891 0.662 3.276 % - -
TPM (Lin et al., 2023) 3.456 0.571 6.929% 0.361 0.734 6.223% 0.850 0.676 5.460% 41.486 0.593
CREAD (Sun et al., 2024) 3.307 0.594 11.236% 0.369 0.738 6.802% 0.865 0.678 5.772% 39.979 0.597
GR (ours) 3.196 0.614 14.981% 0.333 0.753 8.972% 0.815 0.691 7.80% 38.528 0.604
@@
A/B test APP Usage Time +0.112% (p-value=0.01)
Average App Usage Per User +0.087%
Video Consumption Time +0.129%
    In a stable video recommendation system, a 0.1% increase is significant.
@@
MAE $\downarrow$ XAUC $\uparrow$ MAE $\downarrow$ XAUC $\uparrow$
@@
4.1.3. Compared Methods
Considering baseline methods compared in prior studies (Sun et al., 2024; Lin et al., 2023), we compare several state-of-the-art methods (Covington et al., 2016; Zhan et al., 2022; Zhao et al., 2024; Sun et al., 2024; Lin et al., 2023) with our GR. More details of the compared methods are provided in the supplementary material.
4.2. Performance Comparison (RQ1)
4.2.1. Offline Evaluation
Tab. 1 shows the comparative results between GR and six baselines across three datasets. GR achieves consistent improvements in both MAE and XAUC metrics. For watch time prediction, GR maintains superior performance with 4.117% MAE reduction and a 1.917% XAUC improvement on CIKM16. On the KuaiRec, it significantly outperforms the second-best method with a 3.356% MAE reduction and 3.367% XAUC lift. As for Indust dataset, GR exhibits a 3.629% relative decrease in MAE and a 1.001% improvement in XAUC compared to the CREAD, which is a notable enhancement on a real-world business dataset. Regarding watch ratio predictions, while all models gain significantly from eliminating duration bias, GR maintains the best performance, boasting a 7.756% MAE reduction and a 2.033% XAUC improvement. The comprehensive improvements in both MAE and XAUC substantiate GR’s superiority. We also conduct experiments with parameter-equivalent models (see supplementary materials) to ensure the performance gains are not solely from increased model parameters.
4.2.2. Online A/B Testing
[…]
4.3. Underlying Reasons Analysis For Performance Gain (RQ2)
We analyze model performance across ground truth (GT) watch time intervals on KuaiRec, where approximately 80% of videos have GT $\leq$10s. By splitting the range of watch time into 2-second long segments, Fig. 5(a) shows that GR significantly outperforms CREAD and TPM for videos with short and medium watch times, with slightly lower performance only in the last ¿10s interval, where it lags behind TPM. For a more intuitive analysis, Fig. 5(b-d) visualizes the distributions of Ground Truth (GT) watch times and the predictions generated by different methods, alongside their means and variances. Notably, the mean predicted watch time of GR closely aligns with the GT mean, whereas those of CREAD and TPM deviate significantly. Regarding variance, GR exhibits the largest spread, while CREAD shows the smallest. This corresponds visually to CREAD’s highly peaked distribution versus GR’s broader and flatter curve, suggesting GR’s capability to generate a more diverse and personalized set of predictions. Furthermore, GR is the only method that accurately predicts when GT is close to 0s, highlighting its flexibility afforded by its ability to output the ¡EOS¿ token in the first step. Fig. 5(c) and (d) also visually confirm our hypothesis that CREAD and TPM tend to overestimate watch times, stemming from their rigid discretization structure where excessively large span values in tail buckets disproportionately amplify prediction errors, especially for videos with shorter watch times. As shown in Fig. 5(d), the prediction distribution of TPM exhibits a notable skew towards higher values, attributable to the model’s tendency (observed during case analysis) to learn probabilities greater than 0.5 at the root node of its tree structure. This can result in an overall overestimation of the predicted outcomes, thereby explaining why GR’s performance is marginally surpassed by TPM in the ¿10s interval. However, given the characteristic long-tail distribution of real-world watch time data, the superior overall performance and distributional fidelity achieved by GR represent a favorable trade-off for this minor discrepancy in the high-value range.